Quinn can refinish hardwood floors four times as fast as Jack. They have to refinish 100 ft^2 of flooring. Working together, Quinn and Jack can finish the job in 3 hours. How long would it take each of them working alone?
Answer by mananth(16946)

let jack do the job in x hours alone
he does 1/x of the job in 1 hour
quinn does the same job in x/4 hours ( since he does the job 4 times faster)
4/x of the job he does in 1 hour
they together take 3 hours to do the job
1/3 of the job in 1 hour
1/x +4/x = 1/3
5/x= 1/3
x=15 hours jacks's time to do the job alone
Quinn does 4 times faster . he does in 1/4 time = 1/4* 15 = 15/4
=3.75 hours
Ananth